# A Majorization-Minimization Gauss-Newton Method for 1-Bit Matrix Completion

### v2

@@ 3 Related Problems
The 1-bit matrix completion problem considered in this work as well as our approach to solving it are related to several other problems, as we outline below.
Logistic PCA
In logistic PCA, the entries of a binary data matrix ${\bm{\mathbf{{Y}}}}\in\{-1,1\}^{m\times n}$ are modeled as independent draws $y_{ij}\sim\text{Bernoulli}(\Phi(\theta_{ij}^{*}))$ for $(i,j)\in[m]\times[n]$, where ${\bm{\mathbf{{\Theta}}}}^{*}\in\mathbb{R}^{m\times n}$ is rank-$r$ (Collins et al., 2001; Schein et al., 2003). Logistic PCA shares the same modeling assumptions with 1-bit matrix completion but differs in two ways. First, as its name suggests, 1-bit matrix completion only observes part of the data matrix, while logistic PCA assumes all matrix entries are observed. Second, logistic PCA and 1-bit matrix completion have somewhat different goals. The main purpose of the former is dimension reduction, while the latter is prediction of the most likely values at the unobserved entries.
Maximum likelihood algorithms (Collins et al., 2001; Schein et al., 2003; De Leeuw, 2006) have been developed for logistic PCA, which share the same objective function (3). These approaches parameterize ${\bm{\mathbf{{\Theta}}}}$ with an explicit low-rank factorization, i.e., ${\bm{\mathbf{{\Theta}}}}={\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}$. In particular, De Leeuw (2006) developed a majorization-minimization (MM) algorithm for logistic PCA, which employed the same majorization of the objective (3) as in our work. At each MM iteration, the corresponding rank-constrained least squares problem was solved using a weighted singular value decomposition (SVD). In principle, his approach may be adapted to the 1-bit matrix completion problem. However, as the simulations in the supplementary materials demonstrate, when the number of observed entries is small, our method is substantially faster.
Binary matrix factorization
The goal of binary matrix factorization (BMF) is to recover a latent low-rank matrix of probabilities from a binary matrix that is typically fully observed (Bingham et al., 2009; Kabán and Bingham, 2008; Lumbreras et al., 2020; Magron and Févotte, 2022). In the BMF problem, the entries are independent draws from a Bernoulli distribution, i.e., $y_{ij}\sim\text{Bernoulli}([{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}]_{ij})$ where ${\bm{\mathbf{{W}}}}\in\mathbb{R}^{m\times r}$ and ${\bm{\mathbf{{H}}}}\in\mathbb{R}^{n\times r}$. That is, BMF assumes an identity link function, i.e., $\Phi(t)=t$. This is in contrast to employing a distribution function $\Phi$ for the link as in 1-bit matrix completion. In BMF, to guarantee that the entries of ${\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}$ are probabilities, additional constraints must be imposed on the factors matrices, e.g., the rows of ${\bm{\mathbf{{W}}}}$ are assumed to be stochastic and the entries of ${\bm{\mathbf{{H}}}}$ binary.
@@
Link prediction in network analysis
In an unweighted undirected network, links between nodes are represented through a binary adjacency matrix. This binary adjacency matrix may only be partially observed. A fundamental task in network analysis is to predict whether a link is present at a corresponding unobserved entry in the adjacency matrix (Liben-Nowell and Kleinberg, 2003). A variety of link prediction methods have been developed, among which the generative model-based methods (Miller et al., 2009; Zhou, 2015a, b) share similarities with 1-bit matrix completion. To model the binary link $y_{ij}$ between nodes $i$ and $j$ for a network consisting of $n$ nodes, they assume $y_{ij}\sim\text{Bernoulli}(\Phi([{\bm{\mathbf{{Z}}}}{\bm{\mathbf{{W}}}}{\bm{\mathbf{{Z}}}}^{\sf T}]_{ij}))$ where ${\bm{\mathbf{{Z}}}}\in\mathbb{R}^{n\times k}$, ${\bm{\mathbf{{W}}}}\in\mathbb{R}^{k\times k}$, and $\Phi$ is a link function. Assumptions on ${\bm{\mathbf{{Z}}}}$, ${\bm{\mathbf{{W}}}}$, and $\Phi$ vary across different methods. Miller et al. (2009) assumed ${\bm{\mathbf{{Z}}}}$ to be a binary latent feature matrix. Zhou (2015a, b) assumed a Bernoulli-Poisson link $\Phi(x)=1-e^{-x}$, which implies that if the binary observation $y_{ij}=0$, then the latent mean parameter $[{\bm{\mathbf{{X}}}}{\bm{\mathbf{{W}}}}{\bm{\mathbf{{Z}}}}^{\sf T}]_{ij}=0$. Thus, a key difference between link prediction and 1-bit matrix completion is that for some observed matrix entries, there is no uncertainty in the underlying latent mean parameter matrix.
@@ 4 Numerical Studies
In addition, the original problem (2) can be solved by generic optimization approaches. Thus, we also compared MMGN to three general schemes: gradient descent (GD) with backtracking line search as in MMGN, limited-memory BFGS (LBFGS), and Manopt (Boumal et al., 2014) for optimization on the Riemannian manifold of fixed-rank matrices. Details on the implementations appear in the supplementary materials.
@@ 4.1 Simulation Settings
Rank estimation. In practice, we need to estimate the rank from the data. In all simulations, we used the following validation approach to estimate the rank. Given a candidate set of $K$ ranks $\{r_{1},\cdots,r_{K}\}$, we randomly split the observations into a training set (80%) and a validation set (20%). For each candidate rank $r_{k}$, we computed an estimate $\hat{{\bm{\mathbf{{\Theta}}}}}_{k}$ using the training set and then calculated the likelihood at $\hat{{\bm{\mathbf{{\Theta}}}}}_{k}$ on the validation set. We selected the candidate rank that produced the greatest likelihood among all $K$ candidates as our estimated rank $r$. Even with a given rank $r$, TraceNorm and MaxNorm could output a matrix with a higher rank. Consequently, we computed a truncated rank-$r$ SVD on their outputs to return a rank-$r$ matrix.
@@
    where $d_{H}^{2}(p,q)=(\sqrt{p}-\sqrt{q})^{2}+(\sqrt{1-p}-\sqrt{1-q})^{2}$ for $p,q\in[0,1]$. The Hellinger distance $d_{H}^{2}(p,q)$ is a standard metric between two probability distributions. The Hellinger distance between two distributions is nonnegative and is zero when the two distributions are identical. It provides a different perspective on estimation accuracy compared to relative error as we will see later at the end of Section 4.3.
@@
Data generation. In our simulations, we considered two types of low-rank underlying matrices: non-spiky and spiky. We generate matrices of each type as follows.
  • Non-spiky. To generate a non-spiky matrix ${\bm{\mathbf{{\Theta}}}}^{*}\in\mathbb{R}^{m\times n}$ of rank $r^{*}$, we constructed ${\bm{\mathbf{{U}}}}^{*}\in\mathbb{R}^{m\times r^{*}}$, ${\bm{\mathbf{{V}}}}^{*}\in\mathbb{R}^{n\times r^{*}}$ with independent and identically distributed (i.i.d.) entries from a uniform distribution on $[-0.5,0.5]$. We set ${\bm{\mathbf{{\Theta}}}}^{*}={\bm{\mathbf{{U}}}}^{*}{\bm{\mathbf{{V}}}}^{\sf*T}$ and scaled it so that $\|{\bm{\mathbf{{\Theta}}}}^{*}\|_{\infty}=1$. This data generation procedure is the same as in Davenport et al. (2014) and Cai and Zhou (2013).
  • Spiky. Here, we constructed ${\bm{\mathbf{{U}}}}^{*}\in\mathbb{R}^{m\times r^{*}}$, ${\bm{\mathbf{{V}}}}^{*}\in\mathbb{R}^{n\times r^{*}}$ with i.i.d. entries from the $t$-distribution with $\nu$ degrees of freedom. We set ${\bm{\mathbf{{\Theta}}}}^{*}={\bm{\mathbf{{U}}}}^{*}{\bm{\mathbf{{V}}}}^{\sf*T}$, without rescaling. A smaller value of $\nu$ yields a heavier-tailed distribution, resulting in a spikier matrix.
For each simulation, we first generated an underlying matrix ${\bm{\mathbf{{\Theta}}}}^{*}$ and then the binary matrix ${\bm{\mathbf{{Y}}}}$ according to model (1). We randomly selected the set of indices $\Omega$ from a uniform distribution with a user-defined observation fraction $\rho$, namely $\rho=\frac{\lvert\Omega\rvert}{mn}$. We ran $20$ replicates for each setting and reported the median of the three performance metrics for each method.
@@ 4.2.1 Varying the noise level
As discussed in Davenport et al. (2014), the noise level $\sigma$ is a crucial parameter in 1-bit matrix completion. As $\sigma$ tends to zero, the problem becomes ill-posed, as two different positive values for $\theta_{ij}^{*}$ yield the same value $y_{ij}=1$. The first simulation shows the sensitivity of different methods to the noise level $\sigma$. We varied $\sigma$ from $10^{-1.25}$ to $10^{0.25}$ over equally spaced values on a log-linear grid. We set the matrix dimension $m=n=1000$ and rank $r^{*}=1$. The generated underlying matrices can be considered non-spiky, as their average spikiness ratio $s({\bm{\mathbf{{\Theta}}}}^{*})$ is $3.02$ with a standard deviation of $0.07$. For each method, we estimated its rank $r$ from $\{1,2,3,4,5\}$ by the validation procedure described in Section 4.1.
Figure 1 shows the median over 20 replicates of the relative errors, Hellinger distances, and runtimes (in seconds) of all methods with an observation fraction $\rho=0.3$. The left panel shows that all methods produced poor relative errors when the noise level is either too low or too high. This corroborates the results in Davenport et al. (2014) and Cai and Zhou (2013). MaxNorm performed slightly better than the other methods when the noise level was extremely low or high. TraceNorm suffered from the highest relative error over a wide range of noise levels. MMGN and the three generic schemes for solving problem (2) achieved comparable errors when the noise level $\sigma\geq 10^{-0.75}$. The middle panel shows that the Hellinger distance of TraceNorm was significantly larger than those of the other methods, except for $\sigma=10^{-1.25}$ where the Hellinger distance of LBFGS was even larger. As seen in the left and middle panels of Figure 1, the relative errors and Hellinger distances disagreed in their assessments of the methods. The two metrics capture different average errors. We will discuss these differences later. The right panel shows that runtimes of all methods decreased as the noise level grew, and MMGN ran the fastest among all compared methods. MaxNorm was more than $20$ times slower than all the other methods over a wide range of noise levels. In light of this, in all subsequent simulations we supplied MaxNorm with the true rank $r^{*}$ instead of estimating it using the validation procedure.
Figure 1: Probit model: Median over 20 replicates of relative error, Hellinger distance, and runtime versus noise level $\sigma$ for a non-spiky underlying matrix of size $m\times n=1000\times 1000$ and rank $r^{*}=1$ at observation fraction $\rho=0.3$.
@@ 4.2.2 Varying the observation fraction
In the second simulation, we investigated the performance of different methods as a function of the observation fraction $\rho\in\{0.2,0.3,\ldots,1\}$. The underlying matrices were the same as in the simulations in Section 4.2.1. We considered the probit model with a noise level $\sigma=1$. We estimated the rank $r$ from $\{1,2,3,4,5\}$ by the validation procedure described in Section 4.1 for all methods except for MaxNorm which was provided with the true rank $r^{*}$.
Figure 2: Probit model ($\sigma=1$): Median over 20 replicates of relative error, Hellinger distance, and runtime versus observation fraction $\rho$ for a non-spiky underlying matrix of size $m\times n=1000\times 1000$ and rank $r^{*}=1$. The relative errors and Hellinger distances of MMGN, LBFGS, GD, and Manopt overlapped in the left and middle panels, respectively.
Figure 2 shows simulation results of all compared methods. As expected, the relative errors as well as the Hellinger distances of all methods decreased as the observation fraction $\rho$ increased. MMGN behaved comparably with MaxNorm and consistently outperformed TraceNorm. Compared to the three generic schemes, GD, LBFGS, and Manopt, MMGN achieved similar estimation performance while running more than five times faster as seen in the right panel of Figure 2. Moreover, MMGN required substantially less time to achieve comparable estimation accuracy with MaxNorm even when MaxNorm enjoyed the advantage of employing the true rank $r^{*}$.
@@ 4.2.3 Varying the matrix dimension
We now compare the performance of different methods under different matrix dimensions. The underlying matrices were square with dimension $n\in\{1000,1500,2000,2500,3000\}$ for simplicity. All underlying matrices had rank $r^{*}=5$. The average spikiness ratios of the generated underlying matrices were $\{4.61,4.81,5.04,5.09,5.19,5.22\}$ for the corresponding values of $n$. We considered the probit model with $\sigma=0.18$ and fixed the observation fraction $\rho=0.8$. For all methods other than MaxNorm which was given the true rank $r^{*}$, we employed the validation approach to estimate the rank $r$ from a candidate set $\{3,4,5,6,7\}$. The results are shown in Figure 3.
@@
Figure 3: Probit model ($\sigma=0.18$): Median over 20 replicates of relative error, Hellinger distance, and runtime versus matrix dimension $n$ for a square non-spiky underlying matrix of rank $r^{*}=5$ with observation fraction $\rho=0.8$.

### v3

@@ 3 Related Problems
The 1-bit matrix completion problem as well as our approach to solving it are related to several other problems, as we outline below.
Logistic PCA.
In logistic PCA, the entries of a binary data matrix ${\bm{\mathbf{{Y}}}}\in\{-1,1\}^{m\times n}$ are modeled as independent draws $y_{ij}\sim\text{Bernoulli}(\Phi(\theta_{ij}^{*}))$ where ${\bm{\mathbf{{\Theta}}}}^{*}\in\mathbb{R}^{m\times n}$ is rank-$r$ (Collins et al., 2001; Schein et al., 2003). Logistic PCA shares the same modeling assumptions with 1-bit matrix completion but differs in two ways. First, as its name suggests, in 1-bit matrix completion, only part of the data matrix is observed, while in logistic PCA, all matrix entries are observed. Second, logistic PCA and 1-bit matrix completion have somewhat different goals. The main purpose of the former is dimension reduction, while the latter is prediction of the most likely values at the unobserved entries.
Maximum likelihood algorithms (Collins et al., 2001; Schein et al., 2003; De Leeuw, 2006) have been developed for logistic PCA, which share the same objective function (3). These approaches parameterize ${\bm{\mathbf{{\Theta}}}}$ with an explicit low-rank factorization, i.e., ${\bm{\mathbf{{\Theta}}}}={\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}$. In particular, De Leeuw (2006) developed an MM algorithm for logistic PCA, which employed the same majorization of the objective (3) as in our work. At each MM iteration, the corresponding rank-constrained least squares problem was solved using a weighted singular value decomposition (SVD). In principle, his approach may be adapted to the 1-bit matrix completion problem. However, as the simulations in the supplementary materials demonstrate, when the number of observed entries is small, our method is substantially faster.
Binary matrix factorization.
The goal of binary matrix factorization (BMF) is to recover a latent low-rank matrix of probabilities from a binary matrix that is typically fully observed (Bingham et al., 2009; Kabán and Bingham, 2008; Lumbreras et al., 2020; Magron and Févotte, 2022). In the BMF problem, the entries are independent draws from a Bernoulli distribution, i.e., $y_{ij}\sim\text{Bernoulli}([{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}]_{ij})$ where ${\bm{\mathbf{{W}}}}\in\mathbb{R}^{m\times r}$ and ${\bm{\mathbf{{H}}}}\in\mathbb{R}^{n\times r}$. That is, BMF assumes an identity link function $\Phi(\theta)=\theta$. This is in contrast to a distribution function $\Phi$ for the link in 1-bit matrix completion. In BMF, to guarantee that the entries of ${\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}$ are probabilities, additional constraints are imposed on the factor matrices, e.g., the rows of ${\bm{\mathbf{{W}}}}$ are assumed to be stochastic and the entries of ${\bm{\mathbf{{H}}}}$ binary.
@@
Link prediction in network analysis.
In an unweighted undirected network, links between nodes are represented by a binary adjacency matrix. This binary adjacency matrix may be partially observed. A fundamental task in network analysis is to predict whether a link is present at a corresponding unobserved entry in the adjacency matrix (Liben-Nowell and Kleinberg, 2003). A variety of link prediction methods have been developed, among which the generative model-based methods (Miller et al., 2009; Zhou, 2015a, b) share similarities with 1-bit matrix completion. To model the binary link $y_{ij}$ between nodes $i$ and $j$ in a network with $n$ nodes, they assume $y_{ij}\sim\text{Bernoulli}(\Phi([{\bm{\mathbf{{Z}}}}{\bm{\mathbf{{W}}}}{\bm{\mathbf{{Z}}}}^{\sf T}]_{ij}))$ where ${\bm{\mathbf{{Z}}}}\in\mathbb{R}^{n\times k}$, ${\bm{\mathbf{{W}}}}\in\mathbb{R}^{k\times k}$, and $\Phi$ is a link function. Assumptions on ${\bm{\mathbf{{Z}}}}$, ${\bm{\mathbf{{W}}}}$, and $\Phi$ vary across different models. Miller et al. (2009) assumed ${\bm{\mathbf{{Z}}}}$ to be a binary latent feature matrix. Zhou (2015a, b) assumed a Bernoulli-Poisson link $\Phi(\theta)=1-e^{-\theta}$, which implies that if the binary observation $y_{ij}=0$, then the latent mean parameter $[{\bm{\mathbf{{X}}}}{\bm{\mathbf{{W}}}}{\bm{\mathbf{{Z}}}}^{\sf T}]_{ij}=0$. Thus, a key difference between link prediction and 1-bit matrix completion is that for some observed matrix entries, there is no uncertainty in the underlying latent mean parameter matrix.
@@ 4 Numerical Studies
In addition, the original problem (2) can be solved by generic optimization approaches. Thus, we also compared MMGN to three general schemes: gradient descent (GD) with backtracking line search as in MMGN, limited-memory BFGS (LBFGS), and Manopt (Boumal et al., 2014) for optimization on the Riemannian manifold of fixed-rank matrices. Details on the implementations appear in the supplementary materials. We also applied GD and LBGS to ridge regularized version of 1-bit matrix completion and report results on these variations in the supplementary materials.
@@ 4.1 Simulation Settings
Data generation. In our simulations, we considered two types of underlying low-rank matrices: non-spiky and spiky. We generated matrices of each type as follows.
  • Non-spiky. To generate a non-spiky matrix ${\bm{\mathbf{{\Theta}}}}^{*}\in\mathbb{R}^{m\times n}$ of rank $r^{*}$, we constructed ${\bm{\mathbf{{U}}}}^{*}\in\mathbb{R}^{m\times r^{*}}$, ${\bm{\mathbf{{V}}}}^{*}\in\mathbb{R}^{n\times r^{*}}$ with independent and identically distributed (i.i.d.) entries from a uniform distribution on $[-0.5,0.5]$. We set ${\bm{\mathbf{{\Theta}}}}^{*}={\bm{\mathbf{{U}}}}^{*}{\bm{\mathbf{{V}}}}^{\sf*T}$ and scaled it so that $\|{\bm{\mathbf{{\Theta}}}}^{*}\|_{\infty}=1$. This data generation procedure was the same as in Davenport et al. (2014) and Cai and Zhou (2013).
  • Spiky. Here, we constructed ${\bm{\mathbf{{U}}}}^{*}\in\mathbb{R}^{m\times r^{*}}$, ${\bm{\mathbf{{V}}}}^{*}\in\mathbb{R}^{n\times r^{*}}$ with i.i.d. entries from the $t$-distribution with $\nu$ degrees of freedom. We set ${\bm{\mathbf{{\Theta}}}}^{*}={\bm{\mathbf{{U}}}}^{*}{\bm{\mathbf{{V}}}}^{\sf*T}$, without rescaling. A smaller value of $\nu$ yielded a heavier-tailed distribution, resulting in a spikier matrix.
For each simulation, we first generated an underlying low-rank matrix ${\bm{\mathbf{{\Theta}}}}^{*}$. Then we randomly selected the set of indices $\Omega$ from a uniform distribution with a user-defined observation fraction $\rho$, namely $\rho=\frac{\lvert\Omega\rvert}{mn}$. Finally, we produced the binary matrix ${\bm{\mathbf{{Y}}}}$ according to the model (1).
Rank estimation. In practice, we need to estimate the rank from the data. In all simulations, we used the following validation approach to estimate the rank. Given a candidate set of $K$ ranks $\{r_{1},\cdots,r_{K}\}$, we randomly split the observations into a training set (80%) and a validation set (20%). For each candidate rank $r_{k}$, we computed an estimate $\hat{{\bm{\mathbf{{\Theta}}}}}_{k}$ using the training set and then calculated the likelihood of $\hat{{\bm{\mathbf{{\Theta}}}}}_{k}$ on the validation set. We selected the candidate rank with the highest likelihood among all $K$ candidates as our estimated rank $r$. Even when given a rank $r$, TraceNorm and MaxNorm could output a matrix with a higher rank. Consequently, we computed a truncated rank-$r$ SVD on their outputs to return a rank-$r$ matrix.
@@
    where $d_{H}^{2}(p,q)=(\sqrt{p}-\sqrt{q})^{2}+(\sqrt{1-p}-\sqrt{1-q})^{2}$ for $p,q\in[0,1]$. The Hellinger distance $d_{H}^{2}(p,q)$ is a standard metric between two probability distributions. It is nonnegative and equals zero when the two distributions are identical. It provides a different perspective on estimation accuracy compared to relative error as we will see later at the end of Section 4.3.
@@
Some comparison methods in our simulation studies had fairly long run times. Consequently, we ran only 20 replicates for each simulation setting. Additionally, performance measures for some of the comparison methods exhibited large variability. Consequently, to better visualize trends, the median as well as the 25th and 75th quantiles of metrics are shown using box plots.
@@ 4.2.1 Varying the noise level
As discussed in Davenport et al. (2014), the noise level $\sigma$ is a crucial parameter in 1-bit matrix completion. As $\sigma$ tends to zero, the problem becomes ill-posed, as two different positive values for $\theta_{ij}^{*}$ yield the same value $y_{ij}=1$. The first simulation explores the sensitivity of different methods to the noise level $\sigma$. We varied $\sigma$ from $10^{-1.25}$ to $10^{0.25}$ over equally spaced values on a log-linear grid. We set the matrix dimension $m=n=1000$ and rank $r^{*}=1$. The generated underlying matrices are non-spiky with an average spikiness ratio of $3.02$ and a standard deviation of $0.07$. For each method, we estimated its rank $r$ from $\{1,2,3,4,5\}$ by the validation procedure described in Section 4.1.
Figure 1 shows the box plots of relative errors, Hellinger distances, and runtimes (in seconds) of all methods with an observation fraction $\rho=0.3$. The left panel shows that when the noise level is either too low or too high, all methods produced poor relative errors. This corroborates the results in Davenport et al. (2014) and Cai and Zhou (2013). MaxNorm performed slightly better than the other methods when the noise level was extremely low or high. TraceNorm suffered from the highest relative error over a wide range of noise levels. MMGN and the three generic schemes for solving problem (2) achieved comparable errors when the noise level $\sigma\geq 10^{-0.75}$. The middle panel shows that the Hellinger distance of TraceNorm was significantly larger than those of the other methods, except for $\sigma=10^{-1.25}$ where the Hellinger distance of LBFGS was even larger. As seen in the left and middle panels of Figure 1, the relative errors and Hellinger distances disagreed in their assessments of the methods. The two metrics capture different errors. We will discuss these differences later. The right panel shows that runtimes of all methods decreased as the noise level grew, and MMGN ran the fastest among all compared methods. MaxNorm was more than $20$ times slower than all the other methods over a wide range of noise levels. In light of this, in all subsequent simulations we supplied MaxNorm with the true rank $r^{*}$ instead of estimating it using the validation procedure.
Figure 1: Probit model: Relative error, Hellinger distance, and runtime versus noise level $\sigma$ for a non-spiky underlying matrix of size $m\times n=1000\times 1000$ and rank $r^{*}=1$ at observation fraction $\rho=0.3$.
@@ 4.2.2 Varying the observation fraction
In the second simulation, we investigated the performance of different methods as a function of the observation fraction $\rho\in\{0.2,0.3,\ldots,1\}$. The underlying matrices were the same as in the simulations in Section 4.2.1. We considered the probit model with a noise level $\sigma=1$. We estimated the rank $r$ from $\{1,2,3,4,5\}$ by the validation procedure described in Section 4.1 for all methods except for MaxNorm which was provided with the true rank.
Figure 2: Probit model ($\sigma=1$): Relative error, Hellinger distance, and runtime versus observation fraction $\rho$ for a non-spiky underlying matrix of size $m\times n=1000\times 1000$ and rank $r^{*}=1$.
The results of this simulation are shown in Figure 2. As expected, the relative errors as well as the Hellinger distances of all methods decreased as the observation fraction $\rho$ increased. Overall, MMGN behaved comparably with MaxNorm and consistently outperformed TraceNorm. Compared to the three generic schemes, GD, LBFGS, and Manopt, MMGN achieved similar estimation performance while running more than five times faster as seen in the right panel of Figure 2. Moreover, MMGN required substantially less time to achieve comparable estimation accuracy with MaxNorm even when MaxNorm enjoyed the advantage of employing the true rank $r^{*}$.
@@ 4.2.3 Varying the matrix dimension
[…]
@@
Figure 3: Probit model ($\sigma=0.18$): Relative error, Hellinger distance, and runtime versus matrix dimension $n$ for a square non-spiky underlying matrix of rank $r^{*}=5$ with observation fraction $\rho=0.8$.